# Quantum Algorithms for the Pathwise Lasso

### v1

@@ 2.2 Concentration bounds
We revise a few notions of probability theory and concentration bounds, starting with the concept of sub-Gaussian variable.
@@
A random variable $Z$ is sub-Gaussian with parameter $\sigma^{2}$ if
$\mathbb{E}\big{[}e^{t(Z-\mathbb{E}[Z])}\big{]}\leq\exp\left(\frac{t^{2}\sigma^{2}}{2}\right)\quad\text{for all}~{}t\in\mathbb{R}.$
@@
$\mathbb{P}[|Z-\mathbb{E}[Z]|>t]\leq 2\exp{\left(-\frac{t^{2}}{2\sigma^{2}}\right)}\quad\text{for all}~{}t\in\mathbb{R}.$
Fact 6.
@@
More generally, one can use Hoeffding’s inequality to bound the sum of bounded independent random variables.
Fact 7 (Hoeffding’s bound).
@@
$\mathbb{P}[|Z-\mathbb{E}[Z]|\geq t]\leq 2\exp\left(-\frac{2t^{2}}{\sum_{i=1}^{n}(b_{i}-a_{i})^{2}}\right)\quad\text{for all}~{}t>0.$
@@
First note that
$|b|-|\widetilde{b}|\leq\epsilon_{b}\implies\frac{1}{|\widetilde{b}|}\leq\frac{1}{|b|-\epsilon_{b}}\leq\frac{2}{|b|}.$
Then
@@ 2.3 Computational model
Our classical computational model is a classical random-access machine. The input to the Lasso problem is a matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ and a vector $\mathbf{y}\in\mathbb{R}^{n}$, which are stored in a classical-readable read-only memory (ROM). For simplicity, reading any entry of $\mathbf{X}$ and $\mathbf{y}$ takes constant time. The classical computer can write bits to a classical-samplable-and-writable memory. More specifically, the classical computer can write a vector $\mathbf{u}\in\mathbb{R}^{m}$ (or parts of it) into a low-overhead samplable data structure in time $O(m)$, which allows it to sample an index $i\in[m]$ from the probability distribution $\mathcal{D}_{\mathbf{u}}$ in time $O(\operatorname{poly}\log{m})$. We will often refer to the ability to sample from $\mathcal{D}_{\mathbf{u}}$ as having sampling access to the vector $\mathbf{u}$. In this work, we shall assume that all classical-samplable-and-writable memories have size $m=O(\operatorname{poly}{n})$. We assume an arithmetic model in which we ignore issues arising from the fixed-point representation of real numbers. All basic arithmetic operations in this model take constant time.
@@
in time $O(\operatorname{poly}\log(nd))$. The classical computer can also write bits to a quantum-readable classical-writable classical memory (QRAM) [GLM08a, GLM08b, ${}^{+}$23], which can be accessed in superposition by the quantum computer. More specifically, the classical computer can write a vector $\mathbf{u}\in\mathbb{R}^{m}$ (or parts of it) into the memory of a QRAM in time $O(m)$, which allows the quantum computer to invoke an oracle $\mathcal{U}_{\mathbf{u}}$ that performs the mapping
@@
in time $O(\operatorname{poly}\log{m})$. We will often refer to the ability to invoke $\mathcal{U}_{\mathbf{u}}$ as having quantum access to the vector $\mathbf{u}$. In this work, we shall assume that all QRAMs have size $m=O(\operatorname{poly}{n})$. We do not necessarily assume that $\mathbf{y}$ is stored in a QROM. Instead, it can be stored in a QRAM in time $O(n)$ by the classical computer.
@@
Kerenidis and Prakash [Pra14, KP17] introduced a classical data structure (sub-sequentially called KP-tree) to store a vector $\mathbf{u}\in\mathbb{R}^{m}$ to enable the efficient preparation of the state
@@
Our quantum algorithms shall commonly build KP-trees of auxiliary vectors in order to prepare the above states.
Fact 9 (KP-tree [Pra14, KP17, CdW23]).
@@ 2.4 Quantum subroutines
In this section, we review some useful quantum subroutines that shall be used in the rest of the paper, starting with the quantum minimum-finding algorithm from Dürr and Høyer [DH96].
Fact 10 (Quantum minimum finding [DH96]).
Given quantum access to a vector $\mathbf{u}\in\mathbb{R}^{m}$, there is a quantum algorithm that finds $\min_{i\in[m]}u_{i}$ with success probability $1-\delta$ in time $\widetilde{O}\left(\sqrt{m}\log\frac{1}{\delta}\right)$.
The above subroutine was later generalised by Chen and de Wolf [CdW23] in the case when one has quantum access to the entries of $\mathbf{u}$ up to some additive error. We note that a similar result, but in a different setting, appeared in [QCR20].
Fact 11 ([CdW23, Theorem 2.4]).
Let $\delta_{1},\delta_{2}\in(0,1)$ such that $\delta_{2}=O\big{(}\delta_{1}^{2}/(m\log(1/\delta_{1}))\big{)}$, $\epsilon>0$, and $\mathbf{u}\in\mathbb{R}^{m}$. Suppose access to a unitary that maps $|k\rangle|\bar{0}\rangle\mapsto|k\rangle|R_{k}\rangle$ such that, for every $k\in[m]$, after measuring the state $|R_{k}\rangle$, with probability at least $1-\delta_{2}$ the first register $r_{k}$ of the measurement outcome satisfies $|r_{k}-u_{k}|\leq\epsilon$. Then there is a quantum algorithm that finds an index $k$ such that $u_{k}\leq\min_{j\in[m]}u_{j}+2\epsilon$ with probability at least $1-\delta_{1}$ and in time $\widetilde{O}(\sqrt{m}\log(1/\delta_{1}))$.
We shall also need the amplitude estimation subroutine from Brassard et al. [BHMT02].
Fact 12 ([BHMT02, Theorem 12]).
Given a natural number $M$ and access to an $(n+1)$-qubit unitary $U$ satisfying
@@
Finally, the following result, which is based on [CdW23, Theorem 3.4], constructs the mapping $|j\rangle|\bar{0}\rangle\mapsto|j\rangle|R_{j}\rangle$ such that $R_{j}$ holds an approximation to $\mathbf{A}_{j}^{\top}\mathbf{u}$ up to an additive error, where $\mathbf{A}\in\mathbb{R}^{n\times d}$ and $\mathbf{u}\in\mathbb{R}^{n}$ can be accessed quantumly via KP-trees.
@@
Given $\mathbf{A}\in\mathbb{R}^{n\times d}$ and $\mathbf{u}\in\mathbb{R}^{n}$, assume they are stored in KP-trees and we have quantum access to their entries. Let $\delta\in(0,1)$ and $\epsilon>0$. Then there is a quantum algorithm that implements the mapping $|j\rangle|\bar{0}\rangle\mapsto|j\rangle|R_{j}\rangle$ such that, for all $j\in[d]$, after measuring the state $|R_{j}\rangle$, with probability at least $1-\delta$, the outcome $r_{j}$ of the first register satisfies
@@
Fix $j\in[d]$. Consider $\mathbf{A}_{j}^{\top}\mathbf{u}=\sum_{i=1}^{n}A_{ij}u_{i}$. Apply the operator from Fact 9 to create
$\sum_{i=1}^{n}\sqrt{\frac{|u_{i}|}{\|\mathbf{u}\|_{1}}}|i\rangle|s_{i}\rangle|\bar{0}\rangle|0\rangle,$ (4)
in time $O(\operatorname{poly}\log{n})$, where $s_{i}\triangleq\operatorname{sign}(u_{i})$. On the other hand, using query access to $\mathbf{A}$ create the mapping $|i\rangle|s_{i}\rangle|\bar{0}\rangle\mapsto|i\rangle|s_{i}\rangle|s_{i}A_{ij}\rangle$ costing time $O(\operatorname{poly}\log(nd))$. Using this operator on the first three registers in Eq. (4), we obtain
@@
$U_{\rm CR^{+}}:|a\rangle|0\rangle\mapsto\begin{cases}|a\rangle(\sqrt{a}|1\rangle+\sqrt{1-a}|0\rangle)\quad&\text{if}~{}a\in(0,1],\\ |a\rangle|0\rangle\quad&\text{otherwise}.\end{cases}$
Applying $U_{\rm CR^{+}}$ on the last two registers in Eq. (5), we obtain
@@
Here, $|{\phi_{1}}\rangle$ and $|{\phi_{0}}\rangle$ are unit vectors. We now use Fact 12 to get an estimate $\widetilde{a}_{+}$ such that
$|\widetilde{a}_{+}-a_{+}|\leq\frac{\epsilon}{2}\sqrt{a_{+}}\implies\left|\|\mathbf{A}_{j}\|_{\infty}\|\mathbf{u}\|_{1}\widetilde{a}_{+}-\sum_{\begin{subarray}{c}i\in[n]\\ s_{i}A_{ij}>0\end{subarray}}A_{ij}u_{i}\right|$ $\leq\frac{\epsilon}{2}\sqrt{\|\mathbf{A}_{j}\|_{\infty}\|\mathbf{u}\|_{1}\sum_{\begin{subarray}{c}i\in[n]\\ s_{i}A_{ij}>0\end{subarray}}A_{ij}u_{i}}$
$\leq\frac{\epsilon}{2}\|\mathbf{A}_{j}\|_{\infty}\|\mathbf{u}\|_{1}$
in time $O(\epsilon^{-1}\log(1/\delta)\operatorname{poly}\log(nd))$ and probability at least $21-\delta/$ 2. Notice that we know $\|\mathbf{u}\|_{1}$. Similarly, we estimate
@@
$\mathbb{P}\left[Z=\|\mathbf{A}_{j}\|_{1}u_{i}\operatorname{sign}(A_{ij})\right]=\frac{|A_{ij}|}{\|\mathbf{A}_{j}\|_{1}}\quad\text{for}~{}i\in[n].$
@@
Sample $q=2\epsilon^{-2}\ln(2/\delta)$ indices $\{i_{1},\dots,i_{q}\}\subseteq[n]$ from the distribution $\mathcal{D}_{\mathbf{A}_{j}}$ by using sampling access to $\mathbf{A}$ and set $Z_{k}=\|\mathbf{A}_{j}\|_{1}u_{i_{k}}\operatorname{sign}(A_{i_{k}j})$, $k\in[q]$. The algorithm outputs $\widehat{Z}=\frac{1}{q}\sum_{k=1}^{q}Z_{k}$ as an estimate for $\mathbf{A}_{j}^{\top}\mathbf{u}$. Since $Z_{1},\dots,Z_{q}$ are i.i.d. copies of $Z$, a Hoeffding’s bound (Fact 7) gives
@@ 3.1 Lasso solution
$\hat{\bm{\beta}}\in\operatorname*{argmin}_{\bm{\beta}\in\mathbb{R}^{d}}\frac{1}{2}\|\mathbf{y}-\mathbf{X}{\bm{\beta}}\|_{2}^{2}+\lambda\|\bm{\beta}\|_{1}\quad\text{for some}~{}\lambda>0,$ (6)
where $\mathbf{X}\in\mathbb{R}^{n\times d}$ is the design matrix and $\mathbf{y}\in\mathbb{R}^{n}$ is a vector of observations. We will cover optimality conditions for the Lasso problem, the general form for its solution, and continuity and uniqueness conditions. Most of these results can be found in [RZ07, MY12, Tib13, ${}^{+}$18]. We start with obtaining the optimality conditions for the Lasso problem.
Fact 15 ([Tib13, Lemma 1]).
@@
[…]
@@
[…]
@@
[…]
@@
Eq. (7) can be obtained by considering the Karush–Kuhn–Tucker conditions, or sub-gradient optimality conditions. More specifically, a vector $\hat{\bm{\beta}}\in\mathbb{R}^{d}$ is a solution to the Lasso problem if and only if (see e.g. [BL06, Proposition 3.1.5])
@@
where $\partial\|\hat{\bm{\beta}}\|_{1}$ denotes the sub-gradient of the $\ell_{1}$-norm at $\hat{\bm{\beta}}$. Eq. (7) then follows from the fact that the sub-gradients $\mathbf{s}\in\mathbb{R}^{d}$ of $\|\hat{\bm{\beta}}\|_{1}$ are vectors such that $s_{i}=\operatorname{sign}(\hat{\beta}_{i})$ if $\hat{\beta}_{i}\neq 0$ and $|s_{i}|\leq 1$ otherwise.
Moving on, first note that $\bm{\eta}$ is a sub-gradient and that the uniqueness of $\mathbf{X}\hat{\bm{\beta}}$ (Fact 15) implies the uniqueness of $\mathcal{A}$ and $\bm{\eta}$. By the definition of the sub-gradients $\mathbf{s}$ in Eq. (7), we know that $\hat{\bm{\beta}}_{\mathcal{I}}=\mathbf{0}$, and therefore the block $\mathcal{A}$ of Eq. (7) can be written as
@@
where we used the identity $\mathbf{A}^{+}(\mathbf{A}^{\top})^{+}\mathbf{A}^{\top}=\mathbf{A}^{+}$. Any $\mathbf{b}\in\operatorname{null}(\mathbf{X}_{\mathcal{A}})$ produces a valid Lasso solution $\hat{\bm{\beta}}_{\mathcal{A}}$ as long as $\hat{\bm{\beta}}_{\mathcal{A}}$ has the correct signs given by $\bm{\eta}_{\mathcal{A}}$, i.e., $\operatorname{sign}(\hat{\beta}_{i})=\eta_{i}$ for $i\in\mathcal{A}$. This restriction can be written as $\hat{\beta}_{i}\eta_{i}\geq 0$, which is Eq. (9). This completes the proof. ∎
We shall abuse the definition of the Karush-Kuhn-Tucker (KKT) conditions and refer to Eq. (7) as the KKT conditions themselves for the Lasso problem. The set
@@
[…]
@@
By Fact 16, any Lasso solution has $\ell_{2}$-norm
@@
It follows from Fact 16 that the Lasso solution is unique if $\operatorname{null}(\mathbf{X}_{\mathcal{A}})=\{\mathbf{0}\}$ and is given by

### v2

@@ 2.2 Concentration bounds
We revise a few notions of probability theory and concentration bounds. Let $\mathcal{N}(\mathbf{0},\mathbf{\Sigma})$ be the multivariate Gaussian distribution with covariance matrix $\mathbf{\Sigma}\succ\mathbf{0}\in\mathbb{R}^{d\times d}$. We say that a matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ is drawn from the $\mathbf{\Sigma}$-Gaussian ensemble if the rows of $\mathbf{X}$ are sampled i.i.d. according to $\mathcal{N}(\mathbf{0},\mathbf{\Sigma})$. We note that $\mathbf{Z}$ is drawn from the $\mathbf{I}$-Gaussian ensemble if and only if $\mathbf{X}=\mathbf{Z}\sqrt{\mathbf{\Sigma}}$ is drawn from the $\mathbf{\Sigma}$-Gaussian ensemble. We shall use the concept of sub-Gaussian variable.
@@
A random single-variable $z$ is sub-Gaussian with parameter $\sigma^{2}$ if $\mathbb{E}[e^{t(z-\mathbb{E}[z])}]\leq\exp\big{(}\frac{t^{2}\sigma^{2}}{2}\big{)}$ for all $t\in\mathbb{R}$.
@@
$\mathbb{P}[z-\mathbb{E}[z]>t]\leq\exp{\left(-\frac{t^{2}}{2\sigma^{2}}\right)}\quad\text{and}\quad\mathbb{P}[|z-\mathbb{E}[z]|>t]\leq 2\exp{\left(-\frac{t^{2}}{2\sigma^{2}}\right)}\quad\text{for all}\leavevmode\nobreak\ t\in\mathbb{R}.$
Fact 7.
@@
More generally, the sum of bounded independent random variables can be bounded via Hoeffding’s inequality.
Fact 8 (Hoeffding’s bound).
@@
$\mathbb{P}[|z-\mathbb{E}[z]|\geq t]\leq 2\exp\left(-\frac{2t^{2}}{\sum_{i=1}^{n}(b_{i}-a_{i})^{2}}\right)\quad\text{for all}\leavevmode\nobreak\ t>0.$
Lévy’s lemma [76, 63, 3, 110, 72] will also be useful.
Fact 9 (Lévy’s lemma, [3, Corollary 4.4.28]).
Let $f:\mathbb{S}^{d}\to\mathbb{R}$ be a function defined on the $d$-dimensional hypersphere $\mathbb{S}^{d}$. Assume $f$ is $K$-Lipschitz, i.e., $|f(\psi)-f(\phi)|\leq K\|\psi-\phi\|$. Then
$\operatorname*{\mathbb{P}}_{\psi\sim\mu_{\rm Haar}}[|f(\psi)-\mathbb{E}[f]|\geq\epsilon]\leq 2\exp\left(-\frac{d\epsilon^{2}}{4K^{2}}\right)\quad\text{for all}\leavevmode\nobreak\ \epsilon>0.$
@@
First note that $|b|-|\widetilde{b}|\leq\epsilon_{b}\implies\frac{1}{|\widetilde{b}|}\leq\frac{1}{|b|-\epsilon_{b}}\leq\frac{2}{|b|}$. Then
@@ 2.3 Computational model
Our classical computational model is a classical random-access machine. The input to the Lasso problem is a matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ and a vector $\mathbf{y}\in\mathbb{R}^{n}$, which are stored in a classical-readable read-only memory (ROM). For simplicity, reading any entry of $\mathbf{X}$ and $\mathbf{y}$ takes constant time. The classical computer can write bits to a classical-samplable-and-writable memory. More specifically, the classical computer can write a vector $\mathbf{u}\in\mathbb{R}^{m}$ (or parts of it) into a low-overhead samplable data structure in time $O(m)$, which allows it to sample an index $i\in[m]$ from the probability distribution $\mathcal{D}_{\mathbf{u}}$ in time $O(\operatorname{poly}\log{m})$. We refer to the ability to sample from $\mathcal{D}_{\mathbf{u}}$ as having sampling access to the vector $\mathbf{u}$. In this work, all classical-samplable-and-writable memories have size $m=O(n)$. We assume an arithmetic model in which we ignore issues arising from the fixed-point representation of real numbers. All basic arithmetic operations in this model take constant time.
@@
in time $O(\operatorname{poly}\log(nd))$. The classical computer can also write bits to a quantum-readable classical-writable classical memory (QRAM) [42, 43, 1], which can be accessed in superposition by the quantum computer. More specifically, the classical computer can write a vector $\mathbf{u}\in\mathbb{R}^{m}$ (or parts of it) into the memory of a QRAM in time $O(m)$, which allows the quantum computer to invoke an oracle $\mathcal{U}_{\mathbf{u}}$ that performs the mapping
@@
in time $O(\operatorname{poly}\log{m})$. We refer to the ability to invoke $\mathcal{U}_{\mathbf{u}}$ as having quantum access to the vector $\mathbf{u}$. In this work, all QRAMs have size $m=O(n)$. We do not assume that $\mathbf{y}$ has been previously stored in a QROM. Instead, it can be stored in a QRAM in time $O(n)$ by the classical computer.
@@
Kerenidis and Prakash [82, 59] introduced a classical data structure (sub-sequentially called KP-tree) to store a vector $\mathbf{u}\in\mathbb{R}^{m}$ to enable the efficient preparation of the state
@@
We shall commonly use KP-trees of auxiliary vectors in order to prepare the above states.
Fact 11 (KP-tree [82, 59, 23]).
@@
Remark 12.
It can be argued that the QROM/QRAM model is too strong of an assumption and it might not be realistic [52]. Nonetheless, we believe it is a natural and fair quantum equivalent to the classical computational model. If we make the assumption that the input can be classically accessed in constant time and/or sampled in poly-logarithmically time, then a quantum computer should also require, in all fairness, a comparable amount of time to quantumly access the input.
@@ 2.4 Quantum subroutines
In this section, we review some useful quantum subroutines that shall be used in the rest of the paper, starting with the quantum minimum-finding algorithm from Dürr and Høyer [25].
Fact 13 (Quantum minimum finding [25]).
Given quantum access to a vector $\mathbf{u}\in\mathbb{R}^{m}$, there is a quantum algorithm that finds $\min_{i\in[m]}u_{i}$ with success probability $1-\delta$ in time $\widetilde{O}(\sqrt{m}\log\frac{1}{\delta})$.
The above subroutine was later generalised by Chen and de Wolf [23] in the case when one has quantum access to the entries of $\mathbf{u}$ up to some additive error. We note that a similar result, but in a different setting, appeared in [4, 85].
Fact 14 ([23, Theorem 2.4]).
Let $\delta_{1},\delta_{2}\in(0,1)$ such that $\delta_{2}=O\big{(}\delta_{1}^{2}/(m\log(1/\delta_{1}))\big{)}$, $\epsilon>0$, and $\mathbf{u}\in\mathbb{R}^{m}$. Suppose access to a unitary that maps $|k\rangle|\bar{0}\rangle\mapsto|k\rangle|R_{k}\rangle$ such that, for every $k\in[m]$, after measuring the state $|R_{k}\rangle$, with probability at least $1-\delta_{2}$ the first register $r_{k}$ of the measurement outcome satisfies $|r_{k}-u_{k}|\leq\epsilon$. Then there is a quantum algorithm that finds an index $k\in[m]$ such that $u_{k}\leq\min_{j\in[m]}u_{j}+2\epsilon$ with probability at least $1-\delta_{1}$ and in time $\widetilde{O}(\sqrt{m}\log(1/\delta_{1}))$.
We shall also need the amplitude estimation subroutine from Brassard et al. [10].
Fact 15 ([10, Theorem 12]).
Given $M\in\mathbb{N}$ and access to an $(n+1)$-qubit unitary $U$ satisfying
@@
Finally, the following result, which is based on [23, Theorem 3.4], constructs the mapping $|j\rangle|\bar{0}\rangle\mapsto|j\rangle|R_{j}\rangle$ such that $R_{j}$ holds an approximation to $\mathbf{A}_{j}^{\top}\mathbf{u}$ up to an additive error, where $\mathbf{A}\in\mathbb{R}^{n\times d}$ and $\mathbf{u}\in\mathbb{R}^{n}$ can be accessed quantumly via KP-trees.
@@
Given $\mathbf{A}\in\mathbb{R}^{n\times d}$ and $\mathbf{u}\in\mathbb{R}^{n}$, assume they are stored in KP-trees and we have quantum access to their entries. Assume that $\|\mathbf{u}\|_{\infty}$ and $\|\mathbf{A}_{j}\|_{\infty}$, $j\in[d]$, are known. Let $\delta\in(0,1)$ and $\epsilon>0$. Then there is a quantum algorithm that implements the mapping $|j\rangle|\bar{0}\rangle\mapsto|j\rangle|R_{j}\rangle$ such that, for all $j\in[d]$, after measuring the state $|R_{j}\rangle$, with probability at least $1-\delta$, the outcome $r_{j}$ of the first register satisfies
@@
Fix $j\in[d]$. We start by approximating $\mathbf{A}_{j}^{\top}\mathbf{u}=\sum_{i=1}^{n}A_{ij}u_{i}$ up to error $\epsilon\|\mathbf{A}_{j}\|_{\infty}\|\mathbf{u}\|_{1}$. Apply the operator from Fact 11 to create
$\sum_{i=1}^{n}\sqrt{\frac{|u_{i}|}{\|\mathbf{u}\|_{1}}}|i\rangle|s_{i}\rangle|\bar{0}\rangle|0\rangle$ (3)
in time $O(\operatorname{poly}\log{n})$, where $s_{i}\triangleq\operatorname{sign}(u_{i})$. On the other hand, using query access to $\mathbf{A}$ create the mapping $|i\rangle|s_{i}\rangle|\bar{0}\rangle\mapsto|i\rangle|s_{i}\rangle|s_{i}A_{ij}\rangle$ costing time $O(\operatorname{poly}\log(nd))$. Using this operator on the first three registers in Equation 3, we obtain
@@
$U_{\rm CR^{+}}:|a\rangle|0\rangle\mapsto\begin{cases}|a\rangle(\sqrt{a}|1\rangle+\sqrt{1-a}|0\rangle)\quad&\text{if}\leavevmode\nobreak\ a\in(0,1],\\ |a\rangle|0\rangle\quad&\text{otherwise}.\end{cases}$
Applying $U_{\rm CR^{+}}$ on the last two registers in Equation 4, we obtain
@@
Here, $|{\phi_{1}}\rangle$ and $|{\phi_{0}}\rangle$ are unit vectors. We now use Fact 15 to get an estimate $\widetilde{a}_{+}$ such that
$|\widetilde{a}_{+}-a_{+}|\leq\frac{\epsilon}{2}\sqrt{a_{+}}\implies\left|\|\mathbf{A}_{j}\|_{\infty}\|\mathbf{u}\|_{1}\widetilde{a}_{+}-\!\sum_{\begin{subarray}{c}i\in[n]\\ s_{i}A_{ij}>0\end{subarray}}\!\!A_{ij}u_{i}\right|$ $\leq\frac{\epsilon}{2}\sqrt{\|\mathbf{A}_{j}\|_{\infty}\|\mathbf{u}\|_{1}\!\sum_{\begin{subarray}{c}i\in[n]\\ s_{i}A_{ij}>0\end{subarray}}\!\!A_{ij}u_{i}}\leq\frac{\epsilon}{2}\|\mathbf{A}_{j}\|_{\infty}\|\mathbf{u}\|_{1}$
in time $O(\epsilon^{-1}\log(1/\delta)\operatorname{poly}\log(nd))$ and probability at least $21-\delta/$ 2. Notice that we know $\|\mathbf{u}\|_{1}$ since it is stored in a KP-tree. Similarly, we estimate
@@
$\mathbb{P}\left[z=\|\mathbf{A}_{j}\|_{1}u_{i}\operatorname{sign}(A_{ij})\right]=\frac{|A_{ij}|}{\|\mathbf{A}_{j}\|_{1}}\quad\text{for}\leavevmode\nobreak\ i\in[n].$
@@
Sample $q=\lceil 2\epsilon^{-2}\ln(2/\delta)\rceil$ indices $\{i_{1},\dots,i_{q}\}\subseteq[n]$ from the distribution $\mathcal{D}_{\mathbf{A}_{j}}$ by using sampling access to $\mathbf{A}$ and set $z_{k}=\|\mathbf{A}_{j}\|_{1}u_{i_{k}}\operatorname{sign}(A_{i_{k}j})$, $k\in[q]$. The algorithm outputs $\widehat{z}=\frac{1}{q}\sum_{k=1}^{q}z_{k}$ as an estimate for $\mathbf{A}_{j}^{\top}\mathbf{u}$. Since $z_{1},\dots,z_{q}$ are i.i.d. copies of $z$, a Hoeffding’s bound (Fact 8) gives
@@ 3.1 Lasso solution
$\hat{\bm{\beta}}\in\operatorname*{argmin}_{\bm{\beta}\in\mathbb{R}^{d}}\frac{1}{2}\|\mathbf{y}-\mathbf{X}{\bm{\beta}}\|_{2}^{2}+\lambda\|\bm{\beta}\|_{1}\quad\text{for some}\leavevmode\nobreak\ \lambda>0,$
where $\mathbf{X}\in\mathbb{R}^{n\times d}$ is the design matrix and $\mathbf{y}\in\mathbb{R}^{n}$ is a vector of observations. We will cover optimality conditions for the Lasso problem, the general form for its solution, and continuity and uniqueness conditions. Most of these results can be found in [91, 78, 104, 92]. We start with obtaining the optimality conditions for the Lasso problem.
Fact 18 ([104, Lemma 1]).
@@
$\mathbf{X}^{\top}(\mathbf{y}-\mathbf{X}\hat{\bm{\beta}})=\lambda\mathbf{s}\quad\text{where}\leavevmode\nobreak\ s_{i}\in\begin{cases}\{\operatorname{sign}(\hat{\beta}_{i})\}&\text{if}\leavevmode\nobreak\ \hat{\beta}_{i}\neq 0,\\ [-1,1]&\text{if}\leavevmode\nobreak\ \hat{\beta}_{i}=0,\\ \end{cases}\quad\text{for}\leavevmode\nobreak\ i\in[d].$ (5)
@@
$\hat{\bm{\beta}}_{\mathcal{I}}=\mathbf{0}\quad\text{and}\quad\hat{\bm{\beta}}_{\mathcal{A}}=\mathbf{X}_{\mathcal{A}}^{+}(\mathbf{y}-\lambda(\mathbf{X}_{\mathcal{A}}^{+})^{\top}\bm{\eta}_{\mathcal{A}})+\mathbf{b},$ (6)
@@
$\mathbf{b}\in\operatorname{null}(\mathbf{X}_{\mathcal{A}})\quad\text{and}\quad\eta_{i}\big{(}[\mathbf{X}_{\mathcal{A}}^{+}(\mathbf{y}-\lambda(\mathbf{X}_{\mathcal{A}}^{+})^{\top}\bm{\eta}_{\mathcal{A}})]_{i}+b_{i}\big{)}\geq 0\leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ i\in\mathcal{A}.$ (7)
@@
Equation 5 can be obtained by considering the Karush–Kuhn–Tucker conditions, or sub-gradient optimality conditions. More specifically, a vector $\hat{\bm{\beta}}\in\mathbb{R}^{d}$ is a solution to the Lasso problem if and only if (see, e.g., [12, Proposition 3.1.5])
@@
where $\partial\|\hat{\bm{\beta}}\|_{1}$ denotes the sub-gradient of the $\ell_{1}$-norm at $\hat{\bm{\beta}}$. Equation 5 then follows from the fact that the sub-gradients $\mathbf{s}\in\mathbb{R}^{d}$ of $\|\hat{\bm{\beta}}\|_{1}$ are vectors such that $s_{i}=\operatorname{sign}(\hat{\beta}_{i})$ if $\hat{\beta}_{i}\neq 0$ and $|s_{i}|\leq 1$ otherwise.
Moving on, first note that $\bm{\eta}$ is a sub-gradient and that the uniqueness of $\mathbf{X}\hat{\bm{\beta}}$ (Fact 18) implies the uniqueness of $\mathcal{A}$ and $\bm{\eta}$. By the definition of the sub-gradients $\mathbf{s}$ in Equation 5, we know that $\hat{\bm{\beta}}_{\mathcal{I}}=\mathbf{0}$, and therefore the block $\mathcal{A}$ of Equation 5 can be written as
@@
where we used the identity $\mathbf{A}^{+}(\mathbf{A}^{\top})^{+}\mathbf{A}^{\top}=\mathbf{A}^{+}$. Any $\mathbf{b}\in\operatorname{null}(\mathbf{X}_{\mathcal{A}})$ produces a valid Lasso solution $\hat{\bm{\beta}}_{\mathcal{A}}$ as long as $\hat{\bm{\beta}}_{\mathcal{A}}$ has the correct signs given by $\bm{\eta}_{\mathcal{A}}$, i.e., $\operatorname{sign}(\hat{\beta}_{i})=\eta_{i}$ for $i\in\mathcal{A}$. This restriction can be written as $\hat{\beta}_{i}\eta_{i}\geq 0$, which is Equation 7. This completes the proof. ∎
We shall abuse the definition of the Karush-Kuhn-Tucker (KKT) conditions and refer to Equation 5 as the KKT conditions themselves for the Lasso problem. The set
@@
$\hat{\bm{\beta}}_{\mathcal{I}}=\mathbf{0}\quad\text{and}\quad\hat{\bm{\beta}}_{\mathcal{A}}=\mathbf{X}_{\mathcal{A}}^{+}(\mathbf{y}-\lambda(\mathbf{X}_{\mathcal{A}}^{+})^{\top}\bm{\eta}_{\mathcal{A}})$
@@
By Fact 19, any Lasso solution has $\ell_{2}$-norm
@@
It follows from Fact 19 that the Lasso solution is unique if $\operatorname{null}(\mathbf{X}_{\mathcal{A}})=\{\mathbf{0}\}$ and is given by